The time =f(x), in seconds, for a free falling object to fall, x ft is modeled by f(x) = square root of x/16. Worker drops hammer from a building and it hits ground after 4 seconds, what is ht. hammer dropped from?
You have helped me before and I really thank you!! I have solved a few on my own, just really stuck this time. Thanks again!!!

Answer by scott8148(6628)

4=sqrt(x/16) ... squaring gives 16=x/16